OpenStudy (anonymous):

Jordan is jumping up and down on a trampoline. He has a mass of 55 kg, and his maximum speed during his first jump is 8 m/s. 1. Does Jordan have kinetic energy during the entire time he is jumping on the trampoline? 2. When during his jumps is Jordan’s kinetic energy the least? 3. When during his jumps is Jordan’s kinetic energy the greatest? 4. Calculate Jordan’s maximum kinetic energy during his first jump. Use the equation KE=1/2mv^2

OpenStudy (anonymous):

i need help with number 4. Plz. fan and medal

OpenStudy (anonymous):

for the first qn, answer is no, because when he reaches the top position on his jump, his KE would be zero and PE would be maximum

OpenStudy (anonymous):

answer for the second qn is when he is at the top position

OpenStudy (anonymous):

now the kinetic energy is greatest at the point when velocity is maximum

OpenStudy (anonymous):

KE= 1/2 mv^2 = 1/2*55*8*8=?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thx so much dude

OpenStudy (anonymous):

@rohitmathew i cant figure out the last question. can u explain it.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

you are not getting the answer?

OpenStudy (shamim):

K.E=(1/2)*mv^2=0.5*55*8^2=?

OpenStudy (anonymous):

@ksuk3322 @Joel_the_boss

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

I need help with number four plz.

OpenStudy (anonymous):

KE = Mass*Velocity^2 All divided by 2

OpenStudy (anonymous):

The formula KE=1/2mv2 for kinetic energy

OpenStudy (anonymous):

i know. but i cant figure out the answer.